# From Text to Multimodality: Exploring the Evolution and Impact of Large Language Models in Medical Practice

### v5

@@ III-B Medical Imaging
MLLMs are rapidly transforming medical imaging by offering potential for significant improvements in diagnosis, treatment planning, and patient care. These models, capable of processing and interpreting both textual and visual data, allow for a more comprehensive understanding of patient conditions. The MISS framework, proposed by Chen et al., treats medical Visual Question Answering (VQA) as a generative task, achieving excellent results with fewer multimodal datasets and demonstrating the advantages of generative models in practical applications [14]. A key strength of MLLMs lies in their ability to analyze medical images in conjunction with textual information such as radiology reports, clinical notes, and patient history. This integration of multimodal data allows for a more holistic and nuanced understanding of a patient’s condition. Yildirim et al. demonstrate the value of this approach in radiology, arguing that integrating multimodal data can lead to a more comprehensive patient assessment [55]. MLLMs can automate the generation of radiology reports, potentially improving efficiency and accuracy while reducing radiologists’ workload [54]. MLLMs also facilitate visual question answering, enabling clinicians to interact with medical images by asking specific questions and receiving relevant information from the model [73].
Despite advantages, several limitations hinder the adoption of MLLMs in medical imaging. One major challenge is the reliance on high-quality, labeled data. Chen et al. address this issue in their work on the MISS framework, proposing solutions for leveraging limited datasets [14]. Interpreting complex medical images may requires specialized knowledge that current MLLMs do not fully possess. Mehandru et al. stress the need for high-fidelity simulations to accurately assess LLM performance in these complex scenarios [73]. Ethics about fairness in image interpretation is also crucial, as these models can perpetuate existing healthcare disparities if not carefully designed and evaluated. Yildirim et al. discuss these considerations in detail, focusing on design requirements for ethical AI use in radiology [55].
@@ III-C Patient Engagement and Communication
Chatbots and Virtual Assistants: One of the most promising applications of MLLMs in patient engagement is chatbots and virtual assistants. Traditional chatbots often rely on rule-based systems or simple ML models, with limited ability to understand complex queries and provide nuanced responses. MLLMs, however, can understand both text and images to create more natural and engaging conversations and result in improved patient experiences [74, 75, 26].
Personalized Health Recommendations: MLLMs can also be used to generate personalized health recommendations by analyzing patient data and medical knowledge. By integrating information from electronic health records, medical literature, and even patient-provided images, MLLMs provide tailored advice on lifestyle changes, medication adherence, etc [17].
Patient Education: Educating patients improves health outcomes, but traditional methods often rely on static materials that may be difficult to understand or hard to be tailored towards individual needs. MLLMs generates personalized educational materials that are interactive, engaging, and easy to comprehend. The MedSumm framework, for example, utilizes LLMs and VLMs to generate detailed summaries of Hindi-English code-mixed medical queries, integrating visual aids to improve comprehension and support personalized medical care [18].
@@ III-D Research and Development
Literature Review and Knowledge Extraction: MLLMs are proving invaluable for navigating and synthesizing the vast and ever-growing body of biomedical literature. For instance, BioLORD-2023 integrates LLMs with knowledge graphs to achieve state-of-the-art performance in semantic textual similarity, concept representation, and named entity linking, enabling researchers to extract meaningful insights from complex medical texts [76, 77]. Similarly, MedMT5 tries to overcome language barriers in medical research by offering a robust, open-source, multilingual model for the medical domain, allowing for broader access to knowledge across different languages [32].
Drug Discovery: While still in the early stages, MLLMs in drug discovery holds potential [78]. These models can analyze complex biological data, such as protein structures and drug interactions, to identify potential drug targets and accelerate the drug development process. By integrating information from various modalities, MLLMs can facilitate a more holistic understanding of disease mechanisms and drug interactions, potentially leading to the discovery of novel therapeutics and personalized medicine approaches.
@@ III-E Administrative Tasks
Automation of Documentation: MLLMs are transforming clinical documentation by automating tasks such as generating radiology reports [49] and transcribing medical conversations [79]. This automation can free up clinicians’ time, allowing them to focus on patient care rather than paperwork. For example, one study explored the use of LLMs to simplify radiological reports for improved patient comprehension, finding that LLMs can effectively create more accessible reports while acknowledging the need for careful validation to mitigate potential inaccuracies [80].
@@ IV-A Data Limitations and Needs
While the potential of MLLMs in healthcare is significant, their development and evaluation are hindered by limitations in data resources. As highlighted by [62], medicine is inherently multimodal, with data spanning text, imaging, genomics, and more. Yet, current research faces several key challenges related to data:
Scarcity of Large-Scale, Multimodal Datasets: Existing biomedical datasets are often limited in size and scope, particularly those incorporating multiple modalities. Some researchers mitigates the lack of datasets with locally-aligned phrase grounding annotations for complex semantic modeling [81], while other researchers often propose new dataset when releasing new models [82]. The lack of large-scale datasets and their restricted size and scope is a major bottleneck for training robust and generalizable MLLMs for diverse medical tasks, especially when considering the need for datasets that reflect real-world clinical scenarios [64, 73].
Lack of Diversity and Representation: Existing datasets often lack diversity in terms of patient demographics, medical conditions, and healthcare settings. This issue is particularly relevant when considering the potential biases introduced [83]. Chen et al. emphasizes the challenges of few-shot learning in predicting rare disease areas due to limited data [13]. The lack of representation results in biased models that perform poorly on underrepresented populations or specific medical conditions. The reliance on single-language data, primarily English, is also a major concern [84, 85, 32]. It is difficult to access large amounts of domain-specific pre-training data for multiple languages, which makes it difficult to resolve linguistic bias [32].
Challenges in Data Acquisition and Annotation: Obtaining high-quality, annotated multimodal data in healthcare is complex and resource-intensive. [86] notes that medical image annotation is costly and time-consuming, while [87] points to the lack of LLMs trained on medical records. This challenge is complicated by the need for expert annotations [88, 61]. The scarcity of medical image-text pairs for pre-training, due to privacy and cost issues, is another major issue [14]. Additionally, ensuring data privacy and obtaining informed consent are critical ethical considerations that require careful attention, particularly when dealing with sensitive medical information [35].
@@ IV-B1 Fostering Effective Interdisciplinary Collaboration
The development of clinically relevant and useful MultiModal Large Language Models (MLLMs) requires bridging the gap between computer science and medicine. This interdisciplinary challenge calls for collaboration among medical professionals, data scientists, ethicists, and policymakers [89, 90, 65]. Such collaboration is essential to foster a shared understanding of both the technical capabilities of LLMs and the specific needs and constraints of the healthcare domain.
As LLMs become more integrated into healthcare workflows, it is crucial to define the roles and responsibilities of various stakeholders [91]. [10] stresses the importance of incentivizing users, developers, providers, and regulators to prepare for the transformative role of LLMs in evidence-based sectors. This preparation includes establishing clear guidelines for accountability and oversight to ensure the safe and ethical use of these powerful tools in healthcare settings.
Clinicians’ expertise is vital in guiding the development and evaluation of MLLMs to ensure they address real-world clinical needs[92, 67] illustrates how integrating an LLM optimized for diagnostic reasoning into a clinical workflow can improve diagnostic accuracy and comprehensiveness. However, further research is needed to explore effective methods for incorporating clinicians’ feedback and domain expertise throughout the model development process. This ongoing collaboration between healthcare professionals and AI developers is key to creating MLLMs that can truly enhance patient care and clinical decision-making [56, 93, 94, 95].
@@ IV-C Enhancing Knowledge Integration
Beyond textual data, integrating domain-specific knowledge from sources like medical ontologies, knowledge graphs, and clinical guidelines is essential for the effectiveness of Large Language Models (LLMs) in complex medical tasks [81]. A Significant challenge in deploying LLMs for healthcare is addressing the issues of hallucinations and bias. LLMs can generate factually incorrect information and perpetuate biases present in their training data, which is particularly concerning in medical contexts. To tackle this problem, [96] introduces Med-HALT, a benchmark and dataset specifically designed to evaluate and mitigate hallucinations in medical LLMs. This tool emphasizes the critical need to address these issues for safer healthcare applications. Additionally, [97] underscores the importance of incorporating diverse real-world data and domain-specific knowledge to reduce factual inaccuracies and improve the model’s grounding in real-world clinical scenarios.
The development of multilingual models represents another crucial area for advancement in medical LLMs. Most LLMs are trained primarily on English data, which limits their accessibility and applicability in diverse linguistic contexts. The potential of bilingual fine-tuned LLMs, such as Taiyi, can achieve superior performance on biomedical NLP tasks compared to general LLMs [98]. However, more research is needed to develop effective methods for creating and evaluating multilingual medical LLMs that can cater to the needs of diverse patient populations [84, 99].
@@ IV-D Ethical and Regulatory Framework
A key issue is the lack of clear guidelines and regulations specifically tailored for the development, deployment, and evaluation of LLMs in healthcare [87, 89]. Existing frameworks for medical AI may not fully address the unique ethical and legal implications of LLMs, especially in the context of multimodality. This gap in comprehensive guidance creates uncertainty for developers, clinicians, and regulators, which could impede responsible innovation and safe implementation.
Bias, fairness, and transparency are critical concerns in the use of LLMs in healthcare. Several studies highlight the potential for bias due to imbalances in training data [10, 13, 30]. This can result in unfair or inaccurate outcomes, particularly for underrepresented or marginalized populations. The lack of transparency in LLM and MLLM training processes and decision-making mechanisms also raises concerns about accountability and trust [89]. Future research should focus on developing robust methods for identifying, quantifying, and mitigating biases, as well as ensuring transparency in their development and deployment.
Patient privacy and data security are important when using LLMs in healthcare, as they involve processing sensitive patient information. Integrating multiple data modalities in MLLMs adds complexity to data management and raises additional privacy concerns [3]. Developing secure data storage, de-identification techniques, and access control mechanisms are crucial areas for future research. These challenges are particularly evident in specific medical fields, such as dentistry, where the use of LLMs requires robust safeguards to protect patient data [33, 100, 101].
Additional research is needed to define the optimal balance between human oversight and LLM autonomy, and establishing robust governance structures for LLMs in healthcare is essential to ensure accountability and public trust. A framework for evaluating LLMs in healthcare, including a governance layer to ensure accountability and public confidence, has been proposed [41]. Clear guidelines and standards are needed for data governance, model development, performance evaluation, bias mitigation, and transparency. A collaborative approach involving developers, clinicians, ethicists, regulators, and patients is vital for establishing trust and promoting the responsible use of LLMs in healthcare [102].
@@ IV-E Technical Advancements Required
Current modality alignment methods, which aim to bridge the semantic gap between different data types like text and images, often struggle to capture the complex relationships and nuances present in medical data. This limitation hinders the ability of MLLMs to integrate information effectively and generate accurate and coherent outputs [103, 7]. Novel approaches are needed to create more robust and nuanced alignment methods that can capture the complex interdependencies between different modalities, ensuring a more holistic understanding of medical data.
@@
The ”black box” nature of large language models, where their internal workings and decision-making processes remain opaque, is a significant challenge for their deployment in high-stakes medical decisions. Clinicians need to understand the rationale behind AI-generated outputs to trust and validate their recommendations. The lack of transparency and interpretability in current LLMs hinders the ability to identify potential biases, errors, or inconsistencies in their reasoning. Further research to understand how LLMs make decisions, particularly in the context of assessing clinical acuity is needed [31]. Developing methods to make LLMs more transparent and interpretable is crucial for ensuring their safe and responsible use in medical applications [104, 105].
@@
Achieving reliable generalization and robustness across diverse medical contexts, patient populations, and languages is crucial for the real-world deployment of MLLMs. Current models often struggle to generalize beyond their training data, which leads to inaccuracies and biases when applied to new populations or scenarios. The study by Zhang et al. demonstrates that while LLMs can effectively analyze data from specific medical specialties, their performance often decreases when applied to other areas [73]. Additional efforts should focus on developing methods to enhance the generalization capabilities of MLLMs, ensuring that they perform consistently and reliably across different medical contexts, diverse patient populations, and various languages.
@@
The large size and computational demands of MLLMs pose a significant barrier to their deployment in resource-constrained settings. Training and deploying these models require substantial computational power, which can be prohibitively expensive [106, 107]. Developing efficient and scalable models that operates on less powerful devices or with reduced computational resources is crucial for making these technologies more accessible and equitable in healthcare.

### v6

@@ III-B Medical Imaging
MLLMs are rapidly transforming medical imaging by offering potential for significant improvements in diagnosis, treatment planning, and patient care. These models, capable of processing and interpreting both textual and visual data, allow for a more comprehensive understanding of patient conditions. The MISS framework, proposed by Chen et al., treats medical Visual Question Answering (VQA) as a generative task, achieving excellent results with fewer multimodal datasets and demonstrating the advantages of generative models in practical applications [Jiawei2024MISS]. A key strength of MLLMs lies in their ability to analyze medical images in conjunction with textual information such as radiology reports, clinical notes, and patient history. This integration of multimodal data allows for a more holistic and nuanced understanding of a patient’s condition. Yildirim et al. demonstrate the value of this approach in radiology, arguing that integrating multimodal data can lead to a more comprehensive patient assessment [Yildirim2024Multimodal]. MLLMs can automate the generation of radiology reports, potentially improving efficiency and accuracy while reducing radiologists’ workload [Bhayana2024Chatbots]. MLLMs also facilitate visual question answering, enabling clinicians to interact with medical images by asking specific questions and receiving relevant information from the model [Mehandru2024Evaluating].
Despite advantages, several limitations hinder the adoption of MLLMs in medical imaging. One major challenge is the reliance on high-quality, labeled data. Chen et al. address this issue in their work on the MISS framework, proposing solutions for leveraging limited datasets [Jiawei2024MISS]. Interpreting complex medical images may requires specialized knowledge that current MLLMs do not fully possess. Mehandru et al. stress the need for high-fidelity simulations to accurately assess LLM performance in these complex scenarios [Mehandru2024Evaluating]. Ethics about fairness in image interpretation is also crucial, as these models can perpetuate existing healthcare disparities if not carefully designed and evaluated. Yildirim et al. discuss these considerations in detail, focusing on design requirements for ethical AI use in radiology [Yildirim2024Multimodal].
@@ III-C Patient Engagement and Communication
Chatbots and Virtual Assistants: One of the most promising applications of MLLMs in patient engagement is chatbots and virtual assistants. Traditional chatbots often rely on rule-based systems or simple ML models, with limited ability to understand complex queries and provide nuanced responses. MLLMs, however, can understand both text and images to create more natural and engaging conversations and result in improved patient experiences [mihalache2024interpretation, sharma2024depression, niu2024ehr].
Personalized Health Recommendations: MLLMs can also be used to generate personalized health recommendations by analyzing patient data and medical knowledge. By integrating information from electronic health records, medical literature, and even patient-provided images, MLLMs provide tailored advice on lifestyle changes, medication adherence, etc [dao2024llm].
Patient Education: Educating patients improves health outcomes, but traditional methods often rely on static materials that may be difficult to understand or hard to be tailored towards individual needs. MLLMs generates personalized educational materials that are interactive, engaging, and easy to comprehend. The MedSumm framework, for example, utilizes LLMs and VLMs to generate detailed summaries of Hindi-English code-mixed medical queries, integrating visual aids to improve comprehension and support personalized medical care [Akash2024MedSumm].
@@ III-D Research and Development
Literature Review and Knowledge Extraction: MLLMs are proving invaluable for navigating and synthesizing the vast and ever-growing body of biomedical literature. For instance, BioLORD-2023 integrates LLMs with knowledge graphs to achieve state-of-the-art performance in semantic textual similarity, concept representation, and named entity linking, enabling researchers to extract meaningful insights from complex medical texts [Francois2023BioLORD, niu2024large]. Similarly, MedMT5 tries to overcome language barriers in medical research by offering a robust, open-source, multilingual model for the medical domain, allowing for broader access to knowledge across different languages [Iker2024MedMT5].
Drug Discovery: While still in the early stages, MLLMs in drug discovery holds potential [steurer2024multimodal]. These models can analyze complex biological data, such as protein structures and drug interactions, to identify potential drug targets and accelerate the drug development process. By integrating information from various modalities, MLLMs can facilitate a more holistic understanding of disease mechanisms and drug interactions, potentially leading to the discovery of novel therapeutics and personalized medicine approaches.
@@ III-E Administrative Tasks
Automation of Documentation: MLLMs are transforming clinical documentation by automating tasks such as generating radiology reports [Nakaura2024impact] and transcribing medical conversations [Ayo2024Sound]. This automation can free up clinicians’ time, allowing them to focus on patient care rather than paperwork. For example, one study explored the use of LLMs to simplify radiological reports for improved patient comprehension, finding that LLMs can effectively create more accessible reports while acknowledging the need for careful validation to mitigate potential inaccuracies [Artsi2024Large].
@@ IV-A Data Limitations and Needs
While the potential of MLLMs in healthcare is significant, their development and evaluation are hindered by limitations in data resources. As highlighted by [Tu2023Towards], medicine is inherently multimodal, with data spanning text, imaging, genomics, and more. Yet, current research faces several key challenges related to data:
Scarcity of Large-Scale, Multimodal Datasets: Existing biomedical datasets are often limited in size and scope, particularly those incorporating multiple modalities. Some researchers mitigates the lack of datasets with locally-aligned phrase grounding annotations for complex semantic modeling [Boecking2022Making], while other researchers often propose new dataset when releasing new models [Sheng2023BiomedCLIP]. The lack of large-scale datasets and their restricted size and scope is a major bottleneck for training robust and generalizable MLLMs for diverse medical tasks, especially when considering the need for datasets that reflect real-world clinical scenarios [Mingyu2024Health, Mehandru2024Evaluating].
Lack of Diversity and Representation: Existing datasets often lack diversity in terms of patient demographics, medical conditions, and healthcare settings. This issue is particularly relevant when considering the potential biases introduced [Schmidgall2024Addressing]. Chen et al. emphasizes the challenges of few-shot learning in predicting rare disease areas due to limited data [Zekai2023Language]. The lack of representation results in biased models that perform poorly on underrepresented populations or specific medical conditions. The reliance on single-language data, primarily English, is also a major concern [Jin2024Better, Qiu2024Towards, Iker2024MedMT5]. It is difficult to access large amounts of domain-specific pre-training data for multiple languages, which makes it difficult to resolve linguistic bias [Iker2024MedMT5].
Challenges in Data Acquisition and Annotation: Obtaining high-quality, annotated multimodal data in healthcare is complex and resource-intensive. [Khare2021MMBERT] notes that medical image annotation is costly and time-consuming, while [Shah2023Creation] points to the lack of LLMs trained on medical records. This challenge is complicated by the need for expert annotations [He2023Survey, Wu2024PMC]. The scarcity of medical image-text pairs for pre-training, due to privacy and cost issues, is another major issue [Jiawei2024MISS]. Additionally, ensuring data privacy and obtaining informed consent are critical ethical considerations that require careful attention, particularly when dealing with sensitive medical information [Ong2024Ethical].
@@ IV-B1 Fostering Effective Interdisciplinary Collaboration
The development of clinically relevant and useful MultiModal Large Language Models (MLLMs) requires bridging the gap between computer science and medicine. This interdisciplinary challenge calls for collaboration among medical professionals, data scientists, ethicists, and policymakers [Clusmann2023future, Bean2023Exploring, Park2024Assessing]. Such collaboration is essential to foster a shared understanding of both the technical capabilities of LLMs and the specific needs and constraints of the healthcare domain.
As LLMs become more integrated into healthcare workflows, it is crucial to define the roles and responsibilities of various stakeholders [Francois2023Self]. [Harrer2023Attention] stresses the importance of incentivizing users, developers, providers, and regulators to prepare for the transformative role of LLMs in evidence-based sectors. This preparation includes establishing clear guidelines for accountability and oversight to ensure the safe and ethical use of these powerful tools in healthcare settings.
Clinicians’ expertise is vital in guiding the development and evaluation of MLLMs to ensure they address real-world clinical needs[D2023Deciphering, Daniel2023Towards] illustrates how integrating an LLM optimized for diagnostic reasoning into a clinical workflow can improve diagnostic accuracy and comprehensiveness. However, further research is needed to explore effective methods for incorporating clinicians’ feedback and domain expertise throughout the model development process. This ongoing collaboration between healthcare professionals and AI developers is key to creating MLLMs that can truly enhance patient care and clinical decision-making [Wenhao2024Multimodal, Malek2023Toward, Armitage2024Large, Briganti2023clinician].
@@ IV-C Enhancing Knowledge Integration
Beyond textual data, integrating domain-specific knowledge from sources like medical ontologies, knowledge graphs, and clinical guidelines is essential for the effectiveness of Large Language Models (LLMs) in complex medical tasks [Boecking2022Making]. A Significant challenge in deploying LLMs for healthcare is addressing the issues of hallucinations and bias. LLMs can generate factually incorrect information and perpetuate biases present in their training data, which is particularly concerning in medical contexts. To tackle this problem, [Logesh2023Med] introduces Med-HALT, a benchmark and dataset specifically designed to evaluate and mitigate hallucinations in medical LLMs. This tool emphasizes the critical need to address these issues for safer healthcare applications. Additionally, [Guangyu2023ClinicalGPT] underscores the importance of incorporating diverse real-world data and domain-specific knowledge to reduce factual inaccuracies and improve the model’s grounding in real-world clinical scenarios.
The development of multilingual models represents another crucial area for advancement in medical LLMs. Most LLMs are trained primarily on English data, which limits their accessibility and applicability in diverse linguistic contexts. The potential of bilingual fine-tuned LLMs, such as Taiyi, can achieve superior performance on biomedical NLP tasks compared to general LLMs [Luo2024Taiyi]. However, more research is needed to develop effective methods for creating and evaluating multilingual medical LLMs that can cater to the needs of diverse patient populations [Jin2024Better, Grote2024paradigm].
@@ IV-D Ethical and Regulatory Framework
A key issue is the lack of clear guidelines and regulations specifically tailored for the development, deployment, and evaluation of LLMs in healthcare [Shah2023Creation, Clusmann2023future]. Existing frameworks for medical AI may not fully address the unique ethical and legal implications of LLMs, especially in the context of multimodality. This gap in comprehensive guidance creates uncertainty for developers, clinicians, and regulators, which could impede responsible innovation and safe implementation.
Bias, fairness, and transparency are critical concerns in the use of LLMs in healthcare. Several studies highlight the potential for bias due to imbalances in training data [Harrer2023Attention, Zekai2023Language, Jiang2024754]. This can result in unfair or inaccurate outcomes, particularly for underrepresented or marginalized populations. The lack of transparency in LLM and MLLM training processes and decision-making mechanisms also raises concerns about accountability and trust [Clusmann2023future]. Future research should focus on developing robust methods for identifying, quantifying, and mitigating biases, as well as ensuring transparency in their development and deployment.
Patient privacy and data security are important when using LLMs in healthcare, as they involve processing sensitive patient information. Integrating multiple data modalities in MLLMs adds complexity to data management and raises additional privacy concerns [Acosta2022Multimodal]. Developing secure data storage, de-identification techniques, and access control mechanisms are crucial areas for future research. These challenges are particularly evident in specific medical fields, such as dentistry, where the use of LLMs requires robust safeguards to protect patient data [Wang2023Accelerating, Huang2023ChatGPT, Hu2024Advancing].
Additional research is needed to define the optimal balance between human oversight and LLM autonomy, and establishing robust governance structures for LLMs in healthcare is essential to ensure accountability and public trust. A framework for evaluating LLMs in healthcare, including a governance layer to ensure accountability and public confidence, has been proposed [Reddy2023Evaluating]. Clear guidelines and standards are needed for data governance, model development, performance evaluation, bias mitigation, and transparency. A collaborative approach involving developers, clinicians, ethicists, regulators, and patients is vital for establishing trust and promoting the responsible use of LLMs in healthcare [Youssef2023Importance].
@@ IV-E Technical Advancements Required
Current modality alignment methods, which aim to bridge the semantic gap between different data types like text and images, often struggle to capture the complex relationships and nuances present in medical data. This limitation hinders the ability of MLLMs to integrate information effectively and generate accurate and coherent outputs [Jun2024Large, Van2024On]. Novel approaches are needed to create more robust and nuanced alignment methods that can capture the complex interdependencies between different modalities, ensuring a more holistic understanding of medical data.
@@
The ”black box” nature of large language models, where their internal workings and decision-making processes remain opaque, is a significant challenge for their deployment in high-stakes medical decisions. Clinicians need to understand the rationale behind AI-generated outputs to trust and validate their recommendations. The lack of transparency and interpretability in current LLMs hinders the ability to identify potential biases, errors, or inconsistencies in their reasoning. Further research to understand how LLMs make decisions, particularly in the context of assessing clinical acuity is needed [Williams2024Use]. Developing methods to make LLMs more transparent and interpretable is crucial for ensuring their safe and responsible use in medical applications [Alsentzer2023Zero, Aleksa2023Interpretable].
@@
Achieving reliable generalization and robustness across diverse medical contexts, patient populations, and languages is crucial for the real-world deployment of MLLMs. Current models often struggle to generalize beyond their training data, which leads to inaccuracies and biases when applied to new populations or scenarios. The study by Zhang et al. demonstrates that while LLMs can effectively analyze data from specific medical specialties, their performance often decreases when applied to other areas [Mehandru2024Evaluating]. Additional efforts should focus on developing methods to enhance the generalization capabilities of MLLMs, ensuring that they perform consistently and reliably across different medical contexts, diverse patient populations, and various languages.
@@
The large size and computational demands of MLLMs pose a significant barrier to their deployment in resource-constrained settings. Training and deploying these models require substantial computational power, which can be prohibitively expensive [Chunyuan2023LLaVA, Bi2024AI]. Developing efficient and scalable models that operates on less powerful devices or with reduced computational resources is crucial for making these technologies more accessible and equitable in healthcare.